Quantitative: In the 1950s liquid hydrogen was considered as a fuel for a new type of aircraft being developed, the SR-71 BlaKEX AIRCA FLOR V ECE CU DEVELOPUERTO REDUCTION GEAR HEAT EXCHANGER DIFFUSER COMPRESSOR * AFTERBURNER NOZZLE TURBINE COMBUSTORlco P = P, To=1120°C P = P. heat exchanger (5) P =P (7) combustor afterburner afterbu, , n compressor Pz/P2 = 5.5 P. =P T, =Diffuser Air drawn into the engine first encounters the diffuser. The purpose of the diffuser is to reduce the velocity of thc.) Determine the volumetric flow rate leaving the compressor and the power required by the compressor. (2 points) CombustorTurbine The turbine provides sufficient mechanical power to drive the compressor and pump (sometimes additional power is extrand leaving the combustor. Similar to the combustor assume that the combustion products have similar properties to air and ne
Answer #1

Sol Velocity of aire entering engin = 700 ma or - 312 2 mesec Mass flow rate of air entering = 154.1 kesen Inlet conditionsb) Applying steady state condition in inlet Tt V.. Te + V 20 260 Since ₂=0 278 + 312.22 - Tz 2x(1005) -) T2 = 326.491 K Pressfor Volumetrii flow rate we have to find density of flow exiting compressor S,- Ps 2200x10 2874 445.5 S₂ - 17.2 kg/m3 & Volumd) Rate of heat transfer SCB. - inCp (Ta-T3) - 154.) x1005 (1120+273-445.5) - 146739798. 7 Joule see (e) Initial temperaturef. Heat gained by hydrogen Heat loss by the air 22971226.2 - Main Spain (Ta - Ts) 148.32 - Tp-Ts Ts = T4-148.32 Ts a 1244.6k19 Work done by turbin - imu Ts) 18504189.5 - 1.81 X 14325 (1144 - Ts) 716.66 = 1144-To To = 430.33K Now for after burner staAgain using steady state egu for the nozzle To = Tp + v T2 =900°C = 1173k J2C (Ti-T) - To=2647.0kk V₂ = 1721.297 m/s Thrust e